# Zero-shot Generalization in Inventory Management: Train, then Estimate and Decide

### v2

@@ 1 Introduction
The digital revolution, characterized by unprecedented availability of historical data and advances in computational capabilities, has led to a paradigm shift in operations management research toward data-driven methodologies (Mišić and Perakis,, 2020). For example, in inventory management, where simple and well-analyzed heuristics are widely adopted, there is a growing interest in developing and applying machine learning (ML) (Qi et al.,, 2023) and deep reinforcement learning (DRL) algorithms (Gijsbrechts et al.,, 2022). This evolution promises more sophisticated solutions to complex inventory practices where heuristics are still the norm (Boute et al.,, 2022).
@@
However, deploying data-driven ML applications in real-life operations involving sequential decision-making presents significant challenges. One major challenge is the unpredictable, dynamic and constantly changing nature of operational environments (Amiri et al.,, 2023; Cheung et al.,, 2023), requiring trained policies to be robust and adaptable to similar yet unseen conditions during deployment (cf. Kirk et al.,, 2023). For instance, global industries frequently encounter shifts in demand patterns and supply uncertainties, highlighting the limitations of relying on swathes of historical data and algorithms trained on them (Gong and Simchi-Levi,, 2024).
@@
A related challenge involves the decision-makers’ (DMs) often incomplete understanding of the problem environment, henceforth referred to as the problem parameters or parameterization, that influence state transitions. Here, the system’s state (cf. Puterman,, 2014) refers to the current status of all relevant variables at a given point in time, and state transitions are the result of actions taken, leading to a new state. (For example, in the lost sales inventory system (cf. Zipkin,, 2008), the state transitions correspond to the changes to on-hand and pipeline inventory and costs incurred in each period when a certain demand occurs and a certain order is placed. The problem parameters influence state transition probabilities and costs, and include the applicable demand and lead time distributions, the holding cost and penalty costs.)
@@
Clearly, the optimal actions in a problem depend on the problem parameters. Thus, for effective deployment of ML/DRL algorithms in real-world settings, it is crucial to estimate and adapt to the true problem parameters. Unpredictable changes, such as shifts in demand distributions and the introduction of new products, can limit the applicability of historical data. Additionally, DMs may face difficulties in accurately observing their current state, for example, due to inaccurate inventory recordings (Bai et al., 2023b, ). Consequently, DMs are often compelled to re-estimate problem parameters based on recent observations, and such estimates must be frequently updated to avoid decision making based on outdated information.
@@
In most applications of ML/DRL to inventory management, policies are trained for each specific set of parameters (or instance) of the decision problem (see e.g. Oroojlooyjadid et al.,, 2021; Gijsbrechts et al.,, 2022; Maichle et al.,, 2024; Wang and Minner,, 2024). However, when deploying such models in a practical setting, the need to remodel and retrain a policy whenever the estimated problem parameters are updated represents a significant computational challenge. This challenge becomes even more pressing when considering that DMs typically need to make decisions across many tasks in each decision period, such as managing replenishments for thousands of products, and that training a single model can take hours (see Gijsbrechts et al.,, 2022).
@@
Firstly, to formalize the problem setting, we propose Super-Markov Decision Processes as a generalization of Markov Decision Processes (MDPs) (Puterman,, 2014; Porteus,, 2002). These latter models are traditionally used to model individual problem instances in inventory management. In contrast, our proposed Super-Markov Decision Process (Super-MDP) encompasses all MDP instances that can potentially arise from specific parameterizations of a class of inventory problems, treating each instance as a particular representation within the broader Super-MDP framework. Thus, Super-MDPs can unify the sequential decision-making process governed by the DM’s ignorance on the applicable problem parameters.
@@
Moreover, we propose the Train, then Estimate and Decide (TED) framework to train GCAs based on the Super-MDP formulation, and to enable their deployment to unknown environments. To train a GCA, we observe that while problem parameters such as true demand and lead time distributions are unknown in typical inventory applications, DMs have domain knowledge that typically allows them to determine a (possibly wide) range in which the true problem parameters are expected to reside. We later call this range probable parameter space. As a result, we can formulate a Super-MDP tailored to the resulting class of problem instances that the DM needs to address. To effectively train a GCA for this Super-MDP, we sample problem instances from that class, and train a policy using DRL to select actions depending on the dynamic problem state as well as the problem parameters, i.e. we learn a single policy that is applicable for all instances in the class. During the Estimate phase, we continuously collect data while simultaneously estimating problem parameters, thus reducing the uncertainty caused by the policies’ incomplete knowledge of the environment. For the Decide phase, we note that a given set of problem parameter estimates yields a specific instance of an inventory problem, and we take decisions for that instance by deploying the trained GCA based on these parameter estimates without retraining (ZSG).
@@
The TED framework builds upon a rich tradition in classical inventory management, where problem specific policies (e.g. base-stock policies, $(s,S)$ policies, capped base-stock policies etc.) are optimized based on estimated problem parameters (e.g. forecasted demand distribution, see Silver et al.,, 2016; Axsäter,, 2015), and the resulting decisions are deployed in practice. Training new DRL policies whenever parameter estimates are updated is prohibitively time-consuming especially in the case of many products, and the TED framework overcomes this limitation by training a single policy (our GCA) for a wide range of parameters, and as such it is readily applicable without retraining after updating parameter estimates, enabling data-driven decision making for complex decision problems.
@@
To demonstrate the effectiveness of our TED framework, we apply it to a class of inventory problems involving periodic review, and further characterized by lost sales, and possibly by cyclic demand patterns and stochastic lead times. For this class of inventory problems, we train a GCA during the Train phase, named Generally Capable Lost Sales Network (GC-LSN). To validate it, GC-LSN is rigorously benchmarked versus the base-stock and capped base-stock policies on a wide range of problems assuming full availability of the parameters of each problem instance, both for GC-LSN and for the benchmarks. We find that GC-LSN consistently outperforms these well-performing benchmarks. Moreover, as part of the TED framework, we test GC-LSN under conditions where demand and lead time distributions are initially unknown and must be estimated. For the Estimate phase, we adopt the non-parametric Kaplan-Meier estimator (Kaplan and Meier,, 1958) to estimate the demand distribution and construct a relative frequency distribution for the lead time distribution. Our experiments show that TED outperforms a range of benchmarks specifically designed for online learning (Lyu et al.,, 2024; Agrawal and Jia,, 2022; Zhang et al.,, 2020). To our knowledge, TED stands uniquely as the only general-purpose algorithmic framework capable of addressing these diverse inventory challenges collectively, particularly when key information such as demand and lead time distributions are initially unknown and when data is limited and censored.
@@
We make the implementation code of TED publicly available, including the trained model weights of our generally capable agent GC-LSN. This can be accessed through our GitHub repository https://github.com/tarkantemizoz/DynaPlex. Our objective in making the code and the trained agent available is twofold: academically, to support the principles of open science and enhance the reproducibility of reinforcement learning algorithms, as emphasized by Henderson et al., (2018); and managerially, to provide a state-of-the-art, ready-to-deploy lost sales inventory policy; practitioners facing similar inventory challenges may find this resource valuable.
@@ 2.1 Zero-shot generalization for reinforcement learning
Following the taxonomy of Kirk et al., (2023), we address the iid ZSG setting, where training and testing environments are drawn from the same distribution, preparing agents for robust performance across diverse yet distributionally similar instances. Our work falls into the category where environment parameters are not directly observable and must be inferred from interactions, as in Sodhani et al., (2022), but differs in both domain and structure: we focus on inventory management problems, where parameter estimation and mapping to problem instances are well-defined and relatively low-dimensional. This focus allows us to develop scalable, high-performing policies through our TED framework and Super-MDP formulation, enabling agents to act effectively across a wide range of parameterized problem instances without retraining.
Within the ZSG-RL literature, a common formalism is the Contextual Markov Decision Process (CMDP) (Hallak et al.,, 2015), where a context variable influences system’s dynamics and rewards. Our Super-MDP formulation differs in two important ways. First, CMDPs are typically used to model a single environment per episode that changes with context, whereas Super-MDPs are designed for settings where DMs must manage many MDP instances simultaneously, such as hundreds of products with heterogeneous demand distributions, lead times, and costs. Each product thus has its own parameters. Second, rather than treating the context as generic features, we exploit domain knowledge to define bounded ranges for these parameters, enabling the construction of a probable parameter space. Each parameter combination defines a distinct MDP instance, and the collection of these instances constitutes the Super-MDP. This population view aligns naturally with inventory management and supports our theoretical and empirical developments.
Despite its importance, ZSG in inventory management remains relatively underexplored. Notable early work bridging this gap includes Batsis and Samothrakis, (2024), Akkerman et al., (2025), van der Haar et al., (2024), and Dehaybe et al., (2024) which suggest potential for developing GCAs in this area. A key distinction of our work from those applications in inventory management is the introduction of the unifying Super-MDP formulation of the problem instances, which enables modeling cases where true problem parameters are initially unknown. This uncertainty necessitates the Estimate phase in our TED framework, allowing for real-time parameter estimation and policy adaptation. In contrast, prior applications assume that all true parameters are known in advance.
@@ 2.2 Online learning for inventory systems
In real-life scenarios, DMs frequently encounter unknown dynamics in demand and supply. As a result, a significant segment of literature has explored online learning strategies tailored to managing lost sales inventory control with non-zero lead times. This body of work has primarily focused on optimizing policy parameters (e.g. base-stock levels) in online learning, and has recently achieved breakthroughs in the form of worst-case regret bounds for parameter optimization within various heuristic policy classes, including results for base-stock policies (Agrawal and Jia,, 2022), capped-base stock policies (Lyu et al.,, 2024), and constant order policies (Chen et al., 2023a, ). Of course, adoption of heuristics limits decision makers to the performance of those heuristics for various cases. For instance, constant order policies perform poorly when the long-run average cost does not remain insensitive to lead times (Bai et al., 2023a, ), while appropriate heuristics for the case of cyclic demand patterns are missing (Gong and Simchi-Levi,, 2024).
Our study introduces a new approach to real-time decision making in which the policy is learned offline and applied without further learning during execution. Indeed, our GC-LSN is trained prior to being confronted with any specific instance: we adopt a unified Super-MDP formulation and the TED framework for training GCAs. Within TED, we propose to subsequently feed the trained policy network with a relatively straightforward parameter estimation procedure (i.e., based on the Kaplan-Meier estimator (Kaplan and Meier,, 1958)), and focus on empirical performance in numerical experiments. Our approach is demonstrated to complement the online learning literature with a method that performs very well empirically under a wide range of assumptions. The resulting general policy provides a unique tool for benchmarking a variety of online inventory control strategies for a broad range of problems, both in cases where there are benchmarks with worst-case performance guarantees available, and in cases where such benchmarks are not yet available.
@@ 2.3 Machine learning in operations management
ML has extensive applications across various fields of operations management, including supply chain management, revenue management, and healthcare (Mišić and Perakis,, 2020). Typically, ML approaches adhere to the Predict-Then-Optimize (PTO) paradigm, where ML models forecast future scenarios, such as demands, which are then used to inform optimization decisions (Elmachtoub and Grigas,, 2021). However, forecasting becomes particularly challenging in environments where the future is highly stochastic and decisions have far-reaching consequences (Sinclair et al.,, 2023).
An alternative to the PTO paradigm is DRL, which seeks to optimize decisions directly (Boute et al.,, 2022). DRL has been effectively applied in various domains, including multi-echelon supply chains (Harsha et al.,, 2025; Oroojlooyjadid et al.,, 2021), lost sales inventory control (Gijsbrechts et al.,, 2022; Temizöz et al.,, 2025), and large-scale online retail operations (Madeka et al.,, 2022; Andaz et al.,, 2023; Liu et al.,, 2023). Nevertheless, the effectiveness of DRL often hinges on the availability of accurate uncertainty distributions (see, e.g., Temizöz et al.,, 2025) or a substantial dataset of historical transactions that are applicable for the task to be learned (see, e.g., Madeka et al.,, 2022). In scenarios where data is limited and/or censored, and in cases where problem parameters are not fully observable, the utility of the previous DRL applications may diminish.
@@ 3 Preliminaries
This section introduces the fundamental concepts and notation necessary to define the problem setting and the TED framework presented in the subsequent sections. We begin with an introduction to MDPs, a common modeling framework for sequential decision-making (Puterman,, 2014).
@@
A policy $\pi^{\mathbf{p}}$ for an MDP $\mathcal{M}^{\mathbf{p}}$ with parameterization $\mathbf{p}$ is defined as a function that maps states to actions, i.e., $\pi^{\mathbf{p}}:\mathcal{S}^{\mathbf{p}}\to\mathcal{A}^{\mathbf{p}}$. The optimal policy $\pi^{\mathbf{p}*}$ for $\mathcal{M}^{\mathbf{p}}$ minimizes the undiscounted average cost over an infinite horizon, which aligns with steady-state evaluation in classical inventory management (see Puterman,, 2014):
@@
where $\mathbf{s}_{t}$ is the state at time $t$ and $\pi^{\mathbf{p}}(\mathbf{s}_{t})$ is the action taken in state $\mathbf{s}_{t}$ according to policy $\pi^{\mathbf{p}}$. The expectation $\mathbb{E}$ accounts for the stochastic processes governing the state transitions. Throughout this paper, finding policy $\pi^{\mathbf{p}*}$ for $\mathcal{M}^{\mathbf{p}}$ is referred to as solving a task with a parameterization $\mathbf{p}$. While the solution methodologies for finding $\pi^{\mathbf{p}*}$ are vast (Puterman,, 2014), we are interested in finding a single policy (a GCA $\pi_{S}$, such that $\pi_{S}:\mathcal{S}\times\mathcal{P}\to\mathcal{A}$) that is robust to changes in $\mathbf{p}$, and especially one that can perform well when $\mathbf{p}$ is unknown.

### v3

@@ 1 Introduction
The digital revolution, characterized by unprecedented availability of historical data and advances in computational capabilities, has led to a paradigm shift in operations management research toward data-driven methodologies (Mišić and Perakis, 2020). For example, in inventory management, where simple and well-analyzed heuristics are widely adopted, there is a growing interest in developing and applying machine learning (ML) (Qi et al., 2023) and deep reinforcement learning (DRL) algorithms (Gijsbrechts et al., 2022). This evolution promises more sophisticated solutions to complex inventory practices where heuristics are still the norm (Boute et al., 2022).
@@
However, deploying data-driven ML applications in real-life operations involving sequential decision-making presents significant challenges. One major challenge is the unpredictable, dynamic and constantly changing nature of operational environments (Amiri et al., 2023; Cheung et al., 2023), requiring trained policies to be robust and adaptable to similar yet unseen conditions during deployment (cf. Kirk et al., 2023). For instance, global industries frequently encounter shifts in demand patterns and supply uncertainties, highlighting the limitations of relying on swathes of historical data and algorithms trained on them (Gong and Simchi-Levi, 2024).
@@
A related challenge involves the decision-makers’ (DMs) often incomplete understanding of the problem environment, henceforth referred to as the problem parameters or parameterization, that influence state transitions. Here, the system’s state (cf. Puterman, 2014) refers to the current status of all relevant variables at a given point in time, and state transitions are the result of actions taken, leading to a new state. (For example, in the lost sales inventory system (cf. Zipkin, 2008), the state transitions correspond to the changes to on-hand and pipeline inventory and costs incurred in each period when a certain demand occurs and a certain order is placed. The problem parameters influence state transition probabilities and costs, and include the applicable demand and lead time distributions, the holding cost and penalty costs.)
@@
Clearly, the optimal actions in a problem depend on the problem parameters. Thus, for effective deployment of ML/DRL algorithms in real-world settings, it is crucial to estimate and adapt to the true problem parameters. Unpredictable changes, such as shifts in demand distributions and the introduction of new products, can limit the applicability of historical data. Additionally, DMs may be constrained in accurately inferring the true parameters when demand is censored by stockouts and only sales data are observed (Huh et al., 2011). Consequently, DMs are often compelled to re-estimate problem parameters based on recent observations, and such estimates must be frequently updated to avoid decision making based on outdated information.
@@
In most applications of ML/DRL to inventory management, policies are trained for each specific set of parameters (or instance) of the decision problem (see e.g. Oroojlooyjadid et al., 2021; Gijsbrechts et al., 2022; Maichle et al., 2024; Wang and Minner, 2024). However, when deploying such models in a practical setting, the need to remodel and retrain a policy whenever the estimated problem parameters are updated represents a significant computational challenge. This challenge becomes even more pressing when considering that DMs typically need to make decisions across many tasks in each decision period, such as managing replenishments for thousands of products, and that training a single model can take hours (see Gijsbrechts et al., 2022).
@@
Firstly, to formalize the problem setting, we introduce a Super-Markov Decision Process (Super-MDP) representation for a class of inventory problems. Classical Markov Decision Processes (MDPs) (Puterman, 2014; Porteus, 2002) are traditionally used to model individual problem instances in inventory management. In contrast, a Super-MDP encompasses all MDP instances that can potentially arise from specific parameterizations of this class of problems, treating each instance as a particular member of a broader population of related MDPs. This population-based representation captures the DM’s sequential decision-making under uncertainty about which problem parameters apply.
@@
Moreover, we propose the Train, then Estimate and Decide (TED) framework to train GCAs based on the Super-MDP formulation and to enable their deployment in unknown environments. To train a GCA, we observe that while problem parameters such as true demand and lead time distributions are unknown in typical inventory applications, DMs possess domain knowledge that typically allows them to determine a (possibly wide) range in which the true problem parameters are expected to reside. We later call this range probable parameter space. As a result, we can formulate a Super-MDP tailored to the resulting class of problem instances that the DM needs to address. To effectively train a GCA for this Super-MDP, we sample problem instances from that class, and train a policy using DRL to select actions depending on the dynamic problem state as well as the problem parameters, i.e. we learn a single policy that is applicable for all instances in the class. During the Estimate phase, we continuously collect data while simultaneously estimating problem parameters, thus reducing the uncertainty caused by the policies’ incomplete knowledge of the environment. For the Decide phase, we note that a given set of problem parameter estimates yields a specific instance of an inventory problem, and we take decisions for that instance by deploying the trained GCA based on these parameter estimates without retraining (ZSG).
@@
The TED framework builds upon a rich tradition in classical inventory management, where problem specific policies (e.g. base-stock policies, $(s,S)$ policies, capped base-stock policies etc.) are optimized based on estimated problem parameters (e.g. forecasted demand distribution, see Silver et al., 2016; Axsäter, 2015), and the resulting decisions are deployed in practice. Training new DRL policies whenever parameter estimates are updated is prohibitively time-consuming especially in the case of many products, and the TED framework overcomes this limitation by training a single policy (our GCA) for a wide range of parameters, and as such it is readily applicable without retraining after updating parameter estimates, enabling data-driven decision making for complex decision problems.
@@
To demonstrate the effectiveness of our TED framework, we apply it to a class of inventory problems involving periodic review, and further characterized by lost sales, and possibly by cyclic demand patterns and stochastic lead times. For this class of inventory problems, we train a GCA during the Train phase, named Generally Capable Lost Sales Network (GC-LSN). To validate it, GC-LSN is rigorously benchmarked versus the base-stock and capped base-stock policies on a wide range of problems assuming full availability of the parameters of each problem instance, both for GC-LSN and for the benchmarks. We find that GC-LSN consistently outperforms these well-performing benchmarks. Moreover, as part of the TED framework, we test GC-LSN under conditions where demand and lead time distributions are initially unknown and must be estimated. For the Estimate phase, we adopt the non-parametric Kaplan-Meier estimator (Kaplan and Meier, 1958) to estimate the demand distribution and construct a relative frequency distribution for the lead time distribution. Our experiments show that TED outperforms a range of benchmarks specifically designed for online learning (Lyu et al., 2024; Agrawal and Jia, 2022; Zhang et al., 2020). To our knowledge, TED stands uniquely as the only general-purpose algorithmic framework capable of addressing these diverse inventory challenges collectively, particularly when key information such as demand and lead time distributions are initially unknown and when data is limited and censored.
@@
We make the implementation code of TED publicly available, including the trained model weights of our generally capable agent GC-LSN. This can be accessed through our GitHub repository [^1] https://github.com/tarkantemizoz/DynaPlex. Our objective in making the code and the trained agent available is twofold: academically, to support the principles of open science and enhance the reproducibility of reinforcement learning algorithms, as emphasized by Henderson et al. (2018); and managerially, to provide a state-of-the-art, ready-to-deploy lost sales inventory policy; practitioners facing similar inventory challenges may find this resource valuable.
@@ 2.1 Zero-shot generalization for reinforcement learning
Following the taxonomy of Kirk et al. (2023), we address the iid ZSG setting, where training and testing environments are drawn from the same distribution, preparing agents for robust performance across diverse yet distributionally similar instances. Our work falls into the category where environment parameters are not directly observable and must be inferred from interactions, as in Sodhani et al. (2022), but differs in both domain and structure: we focus on inventory management problems, where parameter estimation and mapping to problem instances are well-defined and relatively low-dimensional. This focus allows us to develop scalable, high-performing policies through our TED framework and Super-MDP formulation, enabling agents to act effectively across a wide range of parameterized problem instances without retraining.
Within the ZSG-RL literature, a common formalism is the Contextual Markov Decision Process (CMDP) (Hallak et al., 2015), in which a context variable induces variation in the system’s dynamics and rewards. Our Super-MDP formulation can be viewed as a CMDP equipped with additional structure tailored to operations settings. In our case, the “context” is an interpretable parameter vector (e.g., demand, lead-time, and cost parameters), and each parameter vector defines a distinct MDP instance. The DM must manage a population of such instances, consisting of hundreds of products with heterogeneous characteristics. Importantly, we later endow the underlying parameter space with bounded ranges using domain knowledge and a task-specific metric, which together yield a “probable” parameter space and support a Lipschitz analysis of zero-shot generalization across instances. We refer to this population-based, metric CMDP view as a Super-MDP; it aligns naturally with inventory management and forms the basis for the theoretical and empirical developments in this paper.
Despite its importance, ZSG in inventory management remains relatively underexplored. Notable early work bridging this gap includes Batsis and Samothrakis (2024), Akkerman et al. (2025), van der Haar et al. (2024), and Dehaybe et al. (2024) which suggest potential for developing GCAs in this area. A key distinction of our work from those applications in inventory management is the introduction of the unifying Super-MDP formulation of the problem instances, which enables modeling cases where true problem parameters are initially unknown. This uncertainty necessitates the Estimate phase in our TED framework, allowing for real-time parameter estimation and policy adaptation. In contrast, prior applications assume that all true parameters are known in advance.
@@ 2.2 Online learning for inventory systems
In real-life scenarios, DMs frequently encounter unknown dynamics in demand and supply. As a result, a significant segment of literature has explored online learning strategies tailored to managing lost sales inventory control with non-zero lead times. This body of work has primarily focused on optimizing policy parameters (e.g. base-stock levels) in online learning, and has recently achieved breakthroughs in the form of worst-case regret bounds for parameter optimization within various heuristic policy classes, including results for base-stock policies (Agrawal and Jia, 2022), capped-base stock policies (Lyu et al., 2024), and constant order policies (Chen et al., 2023a). Of course, adoption of heuristics limits decision makers to the performance of those heuristics for various cases. For instance, constant order policies perform poorly when the long-run average cost does not remain insensitive to lead times (Bai et al., 2023), while appropriate heuristics for the case of cyclic demand patterns are missing (Gong and Simchi-Levi, 2024).
Our study introduces a new approach to real-time decision making in which the policy is learned offline and applied without further learning during execution. Indeed, our GC-LSN is trained prior to being confronted with any specific instance: we adopt a unified Super-MDP formulation and the TED framework for training GCAs. Within TED, we propose to subsequently feed the trained policy network with a relatively straightforward parameter estimation procedure (i.e., based on the Kaplan-Meier estimator (Kaplan and Meier, 1958)), and focus on empirical performance in numerical experiments. Our approach is demonstrated to complement the online learning literature with a method that performs very well empirically under a wide range of assumptions. The resulting general policy provides a unique tool for benchmarking a variety of online inventory control strategies for a broad range of problems, both in cases where there are benchmarks with worst-case performance guarantees available, and in cases where such benchmarks are not yet available.
@@ 2.3 Machine learning in operations management
ML has extensive applications across various fields of operations management, including supply chain management, revenue management, and healthcare (Mišić and Perakis, 2020). Typically, ML approaches adhere to the Predict-Then-Optimize (PTO) paradigm, where ML models forecast future scenarios, such as customer demands, which are then used to inform optimization decisions (Elmachtoub and Grigas, 2021). However, forecasting becomes particularly challenging in environments where the future is highly stochastic and decisions have far-reaching consequences (Sinclair et al., 2023).
An alternative to the PTO paradigm is DRL, which seeks to optimize decisions directly (Boute et al., 2022). DRL has been effectively applied in various domains, including multi-echelon supply chains (Harsha et al., 2025; Oroojlooyjadid et al., 2021), lost sales inventory control (Gijsbrechts et al., 2022; Temizöz et al., 2025), and large-scale online retail operations (Madeka et al., 2022; Andaz et al., 2023; Liu et al., 2023). Nevertheless, the effectiveness of DRL often hinges on the availability of accurate uncertainty distributions (see, e.g., Temizöz et al., 2025) or a substantial dataset of historical transactions that are applicable for the task to be learned (see, e.g., Madeka et al., 2022). In scenarios where data is limited and/or censored, and in cases where problem parameters are not fully observable, the utility of the previous DRL applications may diminish.
@@ 3 Preliminaries
This section introduces the fundamental concepts and notation necessary to define the problem setting and the TED framework presented in the subsequent sections. We begin with an introduction to MDPs, a common modeling framework for sequential decision-making (Puterman, 2014).
@@
A policy $\pi^{\mathbf{p}}$ for an MDP $\mathcal{M}^{\mathbf{p}}$ with parameterization $\mathbf{p}$ is defined as a function that maps states to actions, i.e., $\pi^{\mathbf{p}}:\mathcal{S}^{\mathbf{p}}\to\mathcal{A}^{\mathbf{p}}$. The optimal policy $\pi^{\mathbf{p}*}$ for $\mathcal{M}^{\mathbf{p}}$ minimizes the undiscounted average cost over an infinite horizon, which aligns with steady-state evaluation in classical inventory management (see Puterman, 2014):
@@
where $\mathbf{s}_{t}$ is the state at time $t$ and $\pi^{\mathbf{p}}(\mathbf{s}_{t})$ is the action taken in state $\mathbf{s}_{t}$ according to policy $\pi^{\mathbf{p}}$. The expectation $\mathbb{E}$ accounts for the stochastic processes governing the state transitions. Throughout this paper, finding policy $\pi^{\mathbf{p}*}$ for $\mathcal{M}^{\mathbf{p}}$ is referred to as solving a task with a parameterization $\mathbf{p}$. While the solution methodologies for finding $\pi^{\mathbf{p}*}$ are vast (Puterman, 2014), we are interested in finding a single policy (a GCA $\pi_{S}$, such that $\pi_{S}:\mathcal{S}\times\mathcal{P}\to\mathcal{A}$) that is robust to changes in $\mathbf{p}$, and especially one that can perform well when $\mathbf{p}$ is unknown.